# Co-Existence of Private 5G Network and Wireless Hospital Systems

### v1

@@ I Introduction
Wireless communications technologies are increasingly embedded in hospital environments, allowing vital services, such as real-time patient telemetry, wireless infusion pumps, robotic surgery, and high-resolution video transmission in operating rooms [who_emc, medical_iot_review]. With the emergence of wireless fidelity (Wi-Fi) 6/6E, Bluetooth Low Energy (BLE), and private fifth-generation (5G) networks, hospitals are transitioning toward intelligent, data-driven infrastructures. Fig. 1 (Generated using AI) illustrates the operation room, patient wards, and corridors of a smart hospital, highlighting the deployment of wireless systems such as private 5G, Wi-Fi, and Bluetooth. However, this evolution presents substantial electromagnetic coexistence challenges, especially in environments with life-critical systems that are prone to radio frequency (RF) interference.
@@
The electromagnetic landscape within hospitals is unique: dense, multipath-rich, and filled with legacy and emerging wireless systems that coexist within shared or adjacent spectrum [coexistence_ieee2021]. Of particular concern is the use of commercial wireless technologies, such as long-term evolution (4G LTE), Wi-Fi, and 5G, in proximity to mission-sensitive devices operating under stringent electromagnetic compatibility (EMC) requirements [iec_emc]. International guidelines, such as those of the International Electrotechnical Commission (IEC 60601-1-2), the International Commission on Non-Ionizing Radiation Protection (ICNIRP), and International Telecommunication Union (ITU)-T K.91, emphasize the need for careful spectrum planning and on site RF exposure assessment before deploying new technologies in clinical environments [ICNIRP2020, itu_k91].
@@
Despite ongoing research in hospital RF monitoring [rf_meas_hospital, spectrum_clinical_itu], most existing studies are limited in either spatial granularity, temporal resolution, or technological scope. Few have addressed coexistence of the 0.4-6.1 GHz frequency band in the context of an operational private 5G deployment, particularly under controlled conditions where regulatory compliance, safe exposure, and system-level interference are simultaneously validated. Furthermore, most previous work has not focused on identifying interference-free spectral zones, or “spectrum islands”, which can be allocated for safe medical data transmission.
@@
This work presents a long-duration spectrum occupancy campaign at the newly constructed Oulu University Hospital (OYS, Oulun Yliopistollinen Sairaala), where a private 5G network in the 3.9-4.1 GHz band has been deployed for clinical-grade services, such as surgical video transmission and telemetry, as part of the Hola 5G project  [hola5goulu]. The study quantifies in-band and adjacent-band RF activity using calibrated high-resolution equipment, identifies spectrally isolated bands suitable for medical wireless systems, evaluates the coexistence of private 5G deployment with adjacent LTE and Wi-Fi services, and validates electromagnetic exposure levels against ICNIRP and WHO limits, confirming compliance and safety. Although the electromagnetic safety related to 5G deployments is well understood by RF professionals, we include an assessment of electromagnetic exposure to provide physicians, hospital administrators, and other nontechnical personnel with a clear visualization of the safety of the clinical environment.
@@ II-A Overview of RF Measurements in Hospitals
Wireless spectrum monitoring in hospital environments presents significant challenges due to the coexistence of diverse medical devices and complex indoor propagation characteristics. Previous studies have primarily focused on characterizing spectrum occupancy and identifying interference sources within limited frequency bands. For instance, Virk et al. [virk2014spectrum] performed day-long measurements in the 2.4 GHz industrial, scientific, and medical (ISM) band, revealing high occupancy from overlapping medical and non-medical transmissions, which underscores the need for robust interference management. Similarly, Mucchi et al. [LorenzoMucchi2014] analyzed spatio-temporal occupancy and modeled aggregate interference, demonstrating persistent utilization in telemetry channels and intermittent activity in others, emphasizing the importance of continuous spectrum awareness.
This work is partially based on the previous studies [virk2014spectrum]. An 84-day survey by Al Kalaa et al. [AlKalaa2017] in the 2.4 GHz Wi-Fi band reported low average occupancy (10%) with sporadic peaks (50%), indicating that long-duration monitoring is essential to capture transient interference events. More recent work, such as Hossain and Gutierrez [rf_meas_hospital], has extended these analyses to broader bands, confirming temporal and spatial variability in hospital spectrum usage. Recent advancements have incorporated cognitive radio techniques to enhance spectrum efficiency in medical settings. Sodagari et al. [sodagari2018technologies] explored cognitive radio-enabled wireless body area networks (WBANs), highlighting opportunities for dynamic spectrum access to mitigate interference while maintaining low power consumption. Overall, while prior research has advanced spectrum occupancy characterization and interference detection, it often remains confined to narrow bands or short durations, leaving gaps in comprehensive, wideband assessments suitable for emerging 5G integrations.
@@ II-B RF Coexistence and Spectrum Management in Hospitals
The integration of medical equipment with commercial wireless technologies in hospital settings poses persistent coexistence challenges. The unlicensed 2.4 GHz band, utilized by Wi-Fi and Bluetooth, is particularly susceptible to congestion, potentially interfering with critical devices, such as patient monitors and infusion pumps. Reports from the Medical Records Institute and Mobile Healthcare Alliance [WirelessPlanning2007] identify radio frequency interference (RFI) as a primary barrier to wireless healthcare adoption, with risks including telemetry disruptions and equipment malfunctions. Case studies have documented interference from nearby cellular devices affecting infusion pumps [WirelessPlanning2007]. To address these issues, hospitals have implemented spectrum management strategies, including segregation of sensitive communications into dedicated bands.
In the United States, the Federal Communications Commission (FCC) Wireless Medical Telemetry Service (WMTS) allocates protected bands (e.g., 608-614 MHz and 1.4 GHz) for interference-free patient monitoring [WMTS_ASHE], coordinated to minimize external and inter-hospital disruptions [WirelessPlanning2007]. However, these dedicated bands offer limited bandwidth and hinder the adoption of advanced networks like private 5G. Alternative approaches suggest shifting devices to less congested 5 GHz Wi-Fi bands; however, emerging unlicensed technologies, such as LTE-U and MuLTEfire, may introduce new contention and interference issues in these shared spectrum environments [Zinno2018Coexistence]. Literature emphasizes proactive frequency planning and inter-system coordination to mitigate coexistence issues [WirelessPlanning2007]. Recent surveys on electromagnetic fields (EMF)-aware network planning [koppel2023survey] highlight the need for regulatory alignment and optimization techniques to balance exposure limits with network performance in dense hospital environments.
@@
Virk et al. [virk2014spectrum] Hospital ISM band occupancy 2.35–2.50 GHz Long-term occupancy in hospital; no 5G or exposure analysis
Mucchi et al. [LorenzoMucchi2014] Hospital ISM band occupancy 2.35–2.70 GHz Spatio-temporal occupancy analysis and aggregate interference modelling
Al Kalaa et al. [AlKalaa2017] Wireless coexistence testing (Bluetooth/Wi-Fi) ISM bands Standards and EMC evaluation for medical device coexistence
Berlet et al. [berlet2022emergency] 5G slicing for emergency ultrasound n/a (mobile/ambulance) Demonstrated remote diagnostics over 5G; no long-term in-hospital RF validation
Sodagari et al. [sodagari2018technologies] Cognitive radio for medical WBANs Various Challenges and technologies for spectrum sharing in body area networks
Abuella et al. [abuella2021hybrid] Hybrid RF/VLC systems RF/VLC bands Survey on topologies, performance, and applications for hybrid networks
Chiaraviglio et al. [chiaraviglio20225g] 5G in healthcare from COVID-19 onward Various Opportunities and challenges for 5G-enabled healthcare applications
@@ II-C Private 5G in Hospitals
Private 5G networks offer transformative potential for hospitals by providing dedicated and secure connectivity for mission-sensitive applications, such as telemedicine, remote surgery, smart healthcare systems, and safe low-latency patient data transmission. Reviews indicate that 5G’s high bandwidth and low latency can enhance remote diagnostics and robotic surgery through real-time and high-resolution video streaming [devi20235g, chiaraviglio20225g]. Pilot projects have demonstrated effective tele-ultrasound over 5G, achieving minimal delays and high-fidelity imaging [Wu2020]. Berlet et al. [berlet2022emergency] showcased 5G slicing for emergency ultrasound in mobile settings, though without extensive in-hospital validation.
Despite these advancements, deploying 5G in dense hospital RF environments presents challenges, including interference with sensitive equipment (e.g., magnetic resonance imaging (MRI) machines, infusion pumps) and ensuring robust coverage amid RF obstacles, like metal walls. The Food and Drug Administration (FDA) mandates wireless coexistence analyses for new medical devices to mitigate safety risks in shared bands [METLabsCoexistence]. Spectrum policy in various countries allocates mid-band frequencies (e.g., 3.7-4.2 GHz) for private 5G, raising concerns about adjacent-band interference [sorensen2022spectrum]. Empirical studies of private 5G in operational hospitals remain limited, with most relying on simulations or small testbeds that overlook real-world RF complexities [qureshi2021service].
@@ II-D Electromagnetic Exposure and Safety Standards
RF exposure is a critical concern in hospitals, where patients and staff are proximate to multiple RF sources, such as cellular devices, Wi-Fi, and Bluetooth devices. International guidelines from ICNIRP and WHO establish safe limits for RF fields across 100 kHz to 300 GHz, protecting against adverse effects [ICNIRP2020]. Compliance aims to prevent tissue heating and other substantiated impacts. Ambient RF measurements in hospitals typically show levels far below ICNIRP thresholds; for example, Wi-Fi and cellular signals yield power densities orders of magnitude under 10 mW limits [Arribas2022]. Bushberg et al. [Bushberg2025] assessed millimeter-wave exposure from small base stations, confirming compliance with safety standards. Introducing private 5G requires scrutiny of exposure from higher frequencies and denser infrastructure.
Few studies quantify exposure in 5G-hospital contexts, as deployments are at an early stage. Prior work focused on lower-power Wi-Fi, Bluetooth, and telemetry systems. Chiaraviglio et al. [chiaraviglio2021health] analyzed 5G health risks from a communications engineering perspective, concluding no substantiated adverse effects at compliant levels. Dense 5G networks may not increase exposure, as demonstrated by Chiaraviglio et al. [chiaraviglio2021dense], due to efficient power distribution. Plets et al. [plets2024assessment] evaluated uplink exposure from 5G devices, finding levels well within limits. This work contributes empirical data on 5G standalone exposure at 3.99 GHz in a hospital, measuring field strengths during peak traffic against ICNIRP thresholds.
@@ II-E Spectrum Isolation and Interference-Free Bands
Identifying interference-free ”spectrum islands” is vital for reliable medical communications. Dedicated allocations, like WMTS bands, provide protected channels for telemetry, minimizing competition with commercial devices [dedicated_wireless_medical]. Beyond regulated bands, dynamic approaches using cognitive radio enable real-time selection of quiet frequencies, thereby reducing interference risks and enhancing the performance of wireless medical systems [sodagari2018technologies]. However, achieving isolation in hospitals is challenging due to pervasive RF devices spanning sub-1 GHz to 5 GHz. Private 5G introduces opportunities: if contained without leakage, it serves as a spectrum island for high-priority traffic. Hybrid RF/Visible Light Communication (VLC) systems offer further isolation, leveraging VLC’s immunity to RF interference for complementary coverage [abuella2021hybrid].
@@ II-F Going Beyond the State-of-the-Art
This work pioneers the evaluation of private 5G networks in a real hospital operational environment, with only a handful of prior projects, such as those in the Franco-German collaboration focusing on time-sensitive networking for tele-services, including robotic surgery and telemedicine [FrancoGerman5GEcosystem], and in Hong Kong for smart hospital applications enabling remote consultations, training, and telemedicine via low-latency 5G technologies [GSMA2022], conducting experimental 5G work. The novelty of our work lies in long-duration (24-hour), high-resolution spectrum measurements across 0.4–6.1 GHz, assessing coexistence between private 5G (3.9–4.1 GHz) and adjacent Wi-Fi/LTE services, and empirically validating RF exposure against ICNIRP/WHO thresholds. Table I summarizes key prior studies on hospital spectrum usage, coexistence, and interference management, contrasting them with this work.
@@ III Measurement Setup
Measurement Instrumentation and Frequency Coverage: The core instrument employed was the Agilent E4446A spectrum analyzer (SA) [keysightE4446A], a heterodyne receiver capable of wideband signal acquisition from 3 Hz to 44 GHz. In this study, the SA was configured to span the 400 MHz to 6.1 GHz range, encompassing critical wireless bands prevalent in hospital environments, including medical telemetry (400–608 MHz), ISM bands (902–928 MHz, 2.4–2.5 GHz, and 5 GHz), private 5G pilot allocations (3.9–4.1 GHz), Wi-Fi 5 (5.725–5.875 GHz) and Wi-Fi 6 (5.925–6.1 GHz).
@@
RF Frontend and Antenna System: RF signals were received via the OmniLOG PRO antenna, offering omnidirectional coverage from 150 MHz to 18 GHz. This antenna was chosen for its spectral flatness, high sensitivity, and robustness in multipath-rich indoor environments characteristic of hospitals. It exhibits vertical linear polarization and a voltage standing wave ratio (VSWR) below 2.5:1, ensuring effective impedance matching over the operational frequencies. [aaronia_omnilogpro]. The antenna-to-analyzer interconnection employed SubMiniature version A (SMA) coaxial cables, renowned for their application in high-frequency and microwave systems due to threaded coupling and 50 $\Omega$ impedance. [moco2024_rf_sma]
Automation and Remote Control: Remote operation and automated measurements were enabled through a MATLAB-based control suite interfacing with the SA over an Ethernet local area network (LAN). This was supported by a TeleWell TW-EAV510AC router configured exclusively in wired mode to mitigate electromagnetic interference (EMI) from wireless transmissions. Static IP addressing ensured robust connectivity between the control laptop and the SA, thereby stabilizing the measurement framework [telewell_products].

### v2

@@ I Introduction
Wireless communications technologies are increasingly embedded in hospital environments, allowing vital services, such as real-time patient telemetry, wireless infusion pumps, robotic surgery, and high-resolution video transmission in operating rooms [39, 36]. With the emergence of wireless fidelity (Wi-Fi) 6/6E, Bluetooth Low Energy (BLE), and private fifth-generation (5G) networks, hospitals are transitioning toward intelligent, data-driven infrastructures. Fig. 1 (Generated using AI) illustrates the operation room, patient wards, and corridors of a smart hospital, highlighting the deployment of wireless systems such as private 5G, Wi-Fi, and Bluetooth. However, this evolution presents substantial electromagnetic coexistence challenges, especially in environments with life-critical systems that are prone to radio frequency (RF) interference.
@@
The electromagnetic landscape within hospitals is unique: dense, multipath-rich, and filled with legacy and emerging wireless systems that coexist within shared or adjacent spectrum [12]. Of particular concern is the use of commercial wireless technologies, such as long-term evolution (4G LTE), Wi-Fi, and 5G, in proximity to mission-sensitive devices operating under stringent electromagnetic compatibility (EMC) requirements [23]. International guidelines, such as those of the International Electrotechnical Commission (IEC 60601-1-2), the International Commission on Non-Ionizing Radiation Protection (ICNIRP), and International Telecommunication Union (ITU)-T K.91, emphasize the need for careful spectrum planning and on site RF exposure assessment before deploying new technologies in clinical environments [24, 27].
@@
Despite ongoing research in hospital RF monitoring [22, 18], most existing studies are limited in either spatial granularity, temporal resolution, or technological scope. Few have addressed coexistence of the 0.4-6.1 GHz frequency band in the context of an operational private 5G deployment, particularly under controlled conditions where regulatory compliance, safe exposure, and system-level interference are simultaneously validated. Furthermore, most previous work has not focused on identifying interference-free spectral zones, or “spectrum islands”, which can be allocated for safe medical data transmission.
@@
This work presents a long-duration spectrum occupancy campaign at the newly constructed Oulu University Hospital (OYS, Oulun Yliopistollinen Sairaala), where a private 5G network in the 3.9-4.1 GHz band has been deployed for clinical-grade services, such as surgical video transmission and telemetry, as part of the Hola 5G project  [21]. The study quantifies in-band and adjacent-band RF activity using calibrated high-resolution equipment, identifies spectrally isolated bands suitable for medical wireless systems, evaluates the coexistence of private 5G deployment with adjacent LTE and Wi-Fi services, and validates electromagnetic exposure levels against ICNIRP and WHO limits, confirming compliance and safety. Although the electromagnetic safety related to 5G deployments is well understood by RF professionals, we include an assessment of electromagnetic exposure to provide physicians, hospital administrators, and other nontechnical personnel with a clear visualization of the safety of the clinical environment.
@@ II-A Overview of RF Measurements in Hospitals
Wireless spectrum monitoring in hospital environments presents significant challenges due to the coexistence of diverse medical devices and complex indoor propagation characteristics. Previous studies have primarily focused on characterizing spectrum occupancy and identifying interference sources within limited frequency bands. For instance, Virk et al. [38] performed day-long measurements in the 2.4 GHz industrial, scientific, and medical (ISM) band, revealing high occupancy from overlapping medical and non-medical transmissions, which underscores the need for robust interference management. Similarly, Mucchi et al. [31] analyzed spatio-temporal occupancy and modeled aggregate interference, demonstrating persistent utilization in telemetry channels and intermittent activity in others, emphasizing the importance of continuous spectrum awareness.
This work is partially based on the previous studies [38]. An 84-day survey by Al Kalaa et al. [7] in the 2.4 GHz Wi-Fi band reported low average occupancy (10%) with sporadic peaks (50%), indicating that long-duration monitoring is essential to capture transient interference events. More recent work, such as Hossain and Gutierrez [22], has extended these analyses to broader bands, confirming temporal and spatial variability in hospital spectrum usage. Recent advancements have incorporated cognitive radio techniques to enhance spectrum efficiency in medical settings. Sodagari et al. [34] explored cognitive radio-enabled wireless body area networks (WBANs), highlighting opportunities for dynamic spectrum access to mitigate interference while maintaining low power consumption. Overall, while prior research has advanced spectrum occupancy characterization and interference detection, it often remains confined to narrow bands or short durations, leaving gaps in comprehensive, wideband assessments suitable for emerging 5G integrations.
@@ II-B RF Coexistence and Spectrum Management in Hospitals
The integration of medical equipment with commercial wireless technologies in hospital settings poses persistent coexistence challenges. The unlicensed 2.4 GHz band, utilized by Wi-Fi and Bluetooth, is particularly susceptible to congestion, potentially interfering with critical devices, such as patient monitors and infusion pumps. Reports from the Medical Records Institute and Mobile Healthcare Alliance [1] identify radio frequency interference (RFI) as a primary barrier to wireless healthcare adoption, with risks including telemetry disruptions and equipment malfunctions. Case studies have documented interference from nearby cellular devices affecting infusion pumps [1]. To address these issues, hospitals have implemented spectrum management strategies, including segregation of sensitive communications into dedicated bands.
In the United States, the Federal Communications Commission (FCC) Wireless Medical Telemetry Service (WMTS) allocates protected bands (e.g., 608-614 MHz and 1.4 GHz) for interference-free patient monitoring [8], coordinated to minimize external and inter-hospital disruptions [1]. However, these dedicated bands offer limited bandwidth and hinder the adoption of advanced networks like private 5G. Alternative approaches suggest shifting devices to less congested 5 GHz Wi-Fi bands; however, emerging unlicensed technologies, such as LTE-U and MuLTEfire, may introduce new contention and interference issues in these shared spectrum environments [41]. Literature emphasizes proactive frequency planning and inter-system coordination to mitigate coexistence issues [1]. Recent surveys on electromagnetic fields (EMF)-aware network planning [28] highlight the need for regulatory alignment and optimization techniques to balance exposure limits with network performance in dense hospital environments.
@@
Virk et al. [38] Hospital ISM band occupancy 2.35–2.50 GHz Long-term occupancy in hospital; no 5G or exposure analysis
Mucchi et al. [31] Hospital ISM band occupancy 2.35–2.70 GHz Spatio-temporal occupancy analysis and aggregate interference modelling
Al Kalaa et al. [7] Wireless coexistence testing (Bluetooth/Wi-Fi) ISM bands Standards and EMC evaluation for medical device coexistence
Berlet et al. [11] 5G slicing for emergency ultrasound n/a (mobile/ambulance) Demonstrated remote diagnostics over 5G; no long-term in-hospital RF validation
Sodagari et al. [34] Cognitive radio for medical WBANs Various Challenges and technologies for spectrum sharing in body area networks
Abuella et al. [4] Hybrid RF/VLC systems RF/VLC bands Survey on topologies, performance, and applications for hybrid networks
Chiaraviglio et al. [16] 5G in healthcare from COVID-19 onward Various Opportunities and challenges for 5G-enabled healthcare applications
@@ II-C Private 5G in Hospitals
Private 5G networks offer transformative potential for hospitals by providing dedicated and secure connectivity for mission-sensitive applications, such as telemedicine, remote surgery, smart healthcare systems, and safe low-latency patient data transmission. Reviews indicate that 5G’s high bandwidth and low latency can enhance remote diagnostics and robotic surgery through real-time and high-resolution video streaming [17, 16]. Pilot projects have demonstrated effective tele-ultrasound over 5G, achieving minimal delays and high-fidelity imaging [40]. Berlet et al. [11] showcased 5G slicing for emergency ultrasound in mobile settings, though without extensive in-hospital validation.
Despite these advancements, deploying 5G in dense hospital RF environments presents challenges, including interference with sensitive equipment (e.g., magnetic resonance imaging (MRI) machines, infusion pumps) and ensuring robust coverage amid RF obstacles, like metal walls. The Food and Drug Administration (FDA) mandates wireless coexistence analyses for new medical devices to mitigate safety risks in shared bands [29]. Spectrum policy in various countries allocates mid-band frequencies (e.g., 3.7-4.2 GHz) for private 5G, raising concerns about adjacent-band interference [35]. Empirical studies of private 5G in operational hospitals remain limited, with most relying on simulations or small testbeds that overlook real-world RF complexities [33].
@@ II-D Electromagnetic Exposure and Safety Standards
RF exposure is a critical concern in hospitals, where patients and staff are proximate to multiple RF sources, such as cellular devices, Wi-Fi, and Bluetooth devices. International guidelines from ICNIRP and WHO establish safe limits for RF fields across 100 kHz to 300 GHz, protecting against adverse effects [24]. Compliance aims to prevent tissue heating and other substantiated impacts. Ambient RF measurements in hospitals typically show levels far below ICNIRP thresholds; for example, Wi-Fi and cellular signals yield power densities orders of magnitude under 10 mW limits [9]. Bushberg et al. [13] assessed millimeter-wave exposure from small base stations, confirming compliance with safety standards. Introducing private 5G requires scrutiny of exposure from higher frequencies and denser infrastructure.
Few studies quantify exposure in 5G-hospital contexts, as deployments are at an early stage. Prior work focused on lower-power Wi-Fi, Bluetooth, and telemetry systems. Chiaraviglio et al. [14] analyzed 5G health risks from a communications engineering perspective, concluding no substantiated adverse effects at compliant levels. Dense 5G networks may not increase exposure, as demonstrated by Chiaraviglio et al. [15], due to efficient power distribution. Plets et al. [32] evaluated uplink exposure from 5G devices, finding levels well within limits. This work contributes empirical data on 5G standalone exposure at 3.99 GHz in a hospital, measuring field strengths during peak traffic against ICNIRP thresholds.
@@ II-E Spectrum Isolation and Interference-Free Bands
Identifying interference-free ”spectrum islands” is vital for reliable medical communications. Dedicated allocations, like WMTS bands, provide protected channels for telemetry, minimizing competition with commercial devices [10]. Beyond regulated bands, dynamic approaches using cognitive radio enable real-time selection of quiet frequencies, thereby reducing interference risks and enhancing the performance of wireless medical systems [34]. However, achieving isolation in hospitals is challenging due to pervasive RF devices spanning sub-1 GHz to 5 GHz. Private 5G introduces opportunities: if contained without leakage, it serves as a spectrum island for high-priority traffic. Hybrid RF/Visible Light Communication (VLC) systems offer further isolation, leveraging VLC’s immunity to RF interference for complementary coverage [4].
@@ II-F Going Beyond the State-of-the-Art
This work pioneers the evaluation of private 5G networks in a real hospital operational environment, with only a handful of prior projects, such as those in the Franco-German collaboration focusing on time-sensitive networking for tele-services, including robotic surgery and telemedicine [2], and in Hong Kong for smart hospital applications enabling remote consultations, training, and telemedicine via low-latency 5G technologies [20], conducting experimental 5G work. The novelty of our work lies in long-duration (24-hour), high-resolution spectrum measurements across 0.4–6.1 GHz, assessing coexistence between private 5G (3.9–4.1 GHz) and adjacent Wi-Fi/LTE services, and empirically validating RF exposure against ICNIRP/WHO thresholds. Table I summarizes key prior studies on hospital spectrum usage, coexistence, and interference management, contrasting them with this work.
@@ III Measurement Setup
Measurement Instrumentation and Frequency Coverage: The core instrument employed was the Agilent E4446A spectrum analyzer (SA) [5], a heterodyne receiver capable of wideband signal acquisition from 3 Hz to 44 GHz. In this study, the SA was configured to span the 400 MHz to 6.1 GHz range, encompassing critical wireless bands prevalent in hospital environments, including medical telemetry (400–608 MHz), ISM bands (902–928 MHz, 2.4–2.5 GHz, and 5 GHz), private 5G pilot allocations (3.9–4.1 GHz), Wi-Fi 5 (5.725–5.875 GHz) and Wi-Fi 6 (5.925–6.1 GHz).
@@
RF Frontend and Antenna System: RF signals were received via the OmniLOG PRO antenna, offering omnidirectional coverage from 150 MHz to 18 GHz. This antenna was chosen for its spectral flatness, high sensitivity, and robustness in multipath-rich indoor environments characteristic of hospitals. It exhibits vertical linear polarization and a voltage standing wave ratio (VSWR) below 2.5:1, ensuring effective impedance matching over the operational frequencies. [3]. The antenna-to-analyzer interconnection employed SubMiniature version A (SMA) coaxial cables, renowned for their application in high-frequency and microwave systems due to threaded coupling and 50 $\Omega$ impedance. [30]
Automation and Remote Control: Remote operation and automated measurements were enabled through a MATLAB-based control suite interfacing with the SA over an Ethernet local area network (LAN). This was supported by a TeleWell TW-EAV510AC router configured exclusively in wired mode to mitigate electromagnetic interference (EMI) from wireless transmissions. Static IP addressing ensured robust connectivity between the control laptop and the SA, thereby stabilizing the measurement framework [37].